A stationary body is lowered on to a massive platform moving horizontally with a speed of 4 ms^-1. The distance the body slides relative to the platform, if the coefficient of friction is 0.2, is(g=10ms^-2)


a)16m     b)8m    c)4m    d)2m

Grade:11

1 Answers

vikas askiitian expert
13 years ago

applying v2 = u2 + 2as

 finally particle comes to rest over the platform due to friction..

(retardation)a = (cofficient of friction)g =0.2*10=2m/s2

   u2 = 2(2)s

    s = u2/4 = 4m